=== 2019/2/19(Tue.) ===

=== Weekend ===

=== Schedule ===
  * Current status as of 1/22: IR and Green reached X end PD.
  * --(1/22: Surface inspection for PR mirrors.)-- Cleaning will be done after 1/31.
  * --(1/23: Surface inspection for SR mirrors.)--
  * --(1/24-25: Cleaning HR of SR3, SRM.)--

  * --(1/23: BS is aligned to Y-end PD by IR.)--
  * --(1/27: ITMY is aligned to REFL by IR.)--
  * --(1/27: ITMX is aligned to REFL by IR.)--
  * --(1/27: Beam position of IR at SR3 is confirmed how far from the center.)--

  * --(1/28-29: SR3 is aligned to SR2 center by IR.)--
  * --(1/28: Alignment of SR2 -> SRM by IR.)--
  * --(1/29: Green beam is injected from SR2 AR surface to Y end PD.)--

  * --(1/31: Removing bellows at PR and SR area.)--
  * --(1/31-2/1: Cleaning HR of PR3 and HR, AR of SR2)--

  * --(2/5: Alignment of SRM to OMMT1 by IR)--
  * --(2/6: Alignment of ETMY to POS by Green)--

  * --(2/8: Alignment of form OMMT to OSTM by IR)--
  * --(2/8: Alignment of form OSTM to OMC by IR)--

  * --(2/14: Closing door for ITM chambers)--
  * --(2/15: Closing door for other central area)--
  * --(2/15: pumping down for ITMs)--
  * --(2/15: pumping down for other central area)--

==== Note ====
 * Vacuum level ITMX and ITMX chambers are stopped around 10e-3 Pa. VAC team checks why.
 * Type A control loop should be closed all the time to avoid drift or something. If somebody works Type A, please let Lucia know.

==== Report of work ====
 * Type B:
  * BS seems to be healthy. Damping was engaged.
  * SR3(Fabian) and SR2(Enzo) will be checked this morning, then SRM will be checked this afternoon.
  * Needs more time to measure TF after tomorrow.

 * Type A
  * ITMX seems to be fine. ITMY is necessary to be moved.
  * ITMY will be aligned this afternoon.

==== Work plan ====
 * Maintenance will be performed tomorrow morning, for replacing ETMY maybe.